# Data-Driven Goal Recognition Design for General Behavioral Agents

### v1

@@ Abstract
Goal recognition design aims to make limited modifications to decision-making environments with the goal of making it easier to infer the goals of agents acting within those environments. Although various research efforts have been made in goal recognition design, existing approaches are computationally demanding and often assume that agents are (near-)optimal in their decision-making. To address these limitations, we introduce a data-driven approach to goal recognition design that can account for agents with general behavioral models. Following existing literature, we use worst-case distinctiveness (wcd) as a measure of the difficulty in inferring the goal of an agent in a decision-making environment. Our approach begins by training a machine learning model to predict the wcd for a given environment and the agent behavior model. We then propose a gradient-based optimization framework that accommodates various constraints to optimize decision-making environments for enhanced goal recognition. Through extensive simulations, we demonstrate that our approach outperforms existing methods in reducing wcd and enhancing runtime efficiency in conventional setups, and it also adapts to scenarios not previously covered in the literature, such as those involving flexible budget constraints, more complex environments, and suboptimal agent behavior. Moreover, we have conducted human-subject experiments which confirm that our method can create environments that facilitate efficient goal recognition from real-world human decision-makers.
@@ 1 Introduction
With the rapid advancement of artificial intelligence (AI), there has been a surge in interest in human-AI collaboration, aiming to synergize human and AI capabilities across various domains such as gaming, e-commerce, healthcare, industrial engineering, and workflow productivity. Designing AI agents to work alongside humans requires these agents to understand and infer human goals and intentions. While there has been abundant research in goal recognition (Ramírez and Geffner, 2010; Sukthankar et al., 2014) that aims to infer human goals by observing their actions, this work focuses on the goal recognition design problem (Keren et al., 2014), where one needs to identify how to modify decision-making environments to enable better goal recognition.
@@
Our work is built on the goal recognition design problem formulated by Keren et al. (2014). In their work, they proposed the worst-case distinctiveness (wcd) metric, defined as the maximum number of decisions an agent can make without revealing its goal, to measure the difficulty of inferring the agent’s goal. They then aimed to modify the decision-making environment, through removing allowable actions from states, to optimize this measure. Since the introduction of this work, there have been several follow-up works that extend the formulation to deal with different settings, such as stochastic settings (Wayllace et al., 2016, 2017) and partial observability (Keren et al., 2016). More discussion can be found in the survey by Keren et al. (2021).
@@
While there has been significant progress in goal recognition design, there are two main limitations in the literature. First, current approaches require evaluating the difficulty of goal recognition, i.e., worst-case distinctiveness (wcd), for a large number of potential modifications to the decision-making environment. This process is computationally demanding and limits the scalability of these approaches, especially in virtual domains where the number of possible environment configurations is large. Second, decision-making agents in goal recognition design are often assumed to be optimal. While there are efforts addressing suboptimal agents (Keren et al., 2015; Wayllace and Yeoh, 2022), these primarily focus on settings where there is a limited number of deviations from optimal decision-making. This assumption of optimality is limiting in scenarios involving human agents, who are known to often systematically deviate from optimal decision-making due to cognitive and informational constraints (Kahneman, 2003; O’Donoghue and Rabin, 1999).
@@
To address these limitations, we propose a framework for data-driven goal recognition design with general agent behavior. We explicitly incorporate models of agent behavior into the optimization framework to better represent behavioral agents. To tackle the computational challenges, our approach leverages data-driven methods for goal recognition design. The core idea involves building a machine learning oracle that predicts the difficulty of goal recognition (e.g., wcd) given a decision-making environment and an agent’s behavioral model. This oracle is trained on datasets generated from simulations, allowing for the evaluation of wcd in any given environment and agent model. Such an approach significantly accelerates the evaluation of wcd during run-time optimization. Once the machine learning oracle is established, we apply gradient-based optimization to minimize wcd, employing Lagrangian relaxation to manage constraints on environment modifications.
@@
To evaluate our data-driven goal recognition design framework, we conducted a series of simulations. Initially, these simulations were carried out in the standard setup in the grid world envrionment assuming optimal agent behavior. We show that our approach outperforms existing baselines in reducing worst-case distinctiveness (wcd) and demonstrates considerably better run-time efficiency. We then conducted additional simulations to showcase that our approach can generalize to settings beyond the capabilities of existing approaches in the literature. These include scenarios involving flexible budget constraints, more complex environments, and suboptimal agent behavior. Lastly, we have conducted human-subject studies demonstrating that our method can be leveraged to design environments that facilitate efficient goal recognition from real-world human decision-makers. The results highlight the potential of our approach to enable more efficient human-AI collaboration.
Contributions.
The main contributions of this work can be summarized as follows.
@@
    We propose a data-driven optimization framework for goal recognition design. This framework comprises a predictive model that estimates the wcd for a given environment and a model of agent behavior. Leveraging the differentiable property of the predictive model, we utilize a gradient-based optimization method for goal recognition design. The framework is flexible to accommodate various environments, design spaces, and agent models.
@@
    Through extensive simulations, we show that our framework not only outperforms existing approaches in goal recognition design within standard settings but also adapts to scenarios that existing approaches cannot handle. These scenarios include those with general optimization criteria, complex environments, and suboptimal agent behavior.
@@
    Through real-world human-subject experiments, we demonstrate that our approach can adapt to agent models trained on human behavior. Furthermore, our framework can create environments that enable more effective goal recognition with real-world human decision-makers.
@@ 2 Related Work
Our work contributes to the expanding field of human-AI collaboration. Recent research has indicated that optimizing AI alone is insufficient for maximizing the performance of human-AI teams (Bansal et al., 2021; Carroll et al., 2019). To develop truly collaborative AI agents, it is crucial to equip them with the ability to comprehend and predict the intentions and goals of their human counterparts. This challenge lies at the heart of goal recognition research (Kautz and others, 1991; Ramírez and Geffner, 2010; Hong, 2001; Sukthankar et al., 2014; Pereira et al., 2017; Masters et al., 2021). Our work specifically focuses on goal recognition design, an extension of goal recognition that includes modifying environments to better facilitate the process of recognizing goals.
Goal recognition design was originally formulated by Keren et al. (2014). Since this seminal work, numerous research efforts have expanded the concept to accommodate stochastic environments (Wayllace et al., 2017; Wayllace and Yeoh, 2022), different levels of observability (Keren et al., 2016; Wayllace et al., 2020), and a variety of design spaces (Mirsky et al., 2019). The studies most closely aligned with our approach focus on suboptimal agents (Keren et al., 2015; Wayllace and Yeoh, 2022). However, these studies characterize suboptimality by limiting deviations from the optimal policy, a method that may not adequately represent behavior of human agents, who frequently deviate systematically from optimal decision-making. Additionally, most existing work requires evaluating the difficulty of goal recognition in run time for numerous environmental modifications, which constrains the scalability of these methods. Our work sets itself apart by broadening the goal recognition design question to incorporate general agent behavioral models and by implementing a data-driven optimization approach.
From a technical standpoint, we adopt a data-driven optimization approach for goal recognition design. The use of data-driven tools in optimization has gained increasing prominence in the field of mechanism design, as evidenced by a range of recent studies (Dütting et al., 2019; Feng et al., 2018; Golowich et al., 2018; Curry et al., 2020; Kuo et al., 2020; Rahme et al., 2020; Peri et al., 2021; Cornelisse et al., 2022; Yu et al., 2023). In this regard, our work closely aligns with the studies by Yu and Ho (2022) and Yu et al. (2023), who explicitly encoded human behavior and responses within their mechanism designs for optimizing the information policy in environment and information design. While our work shares a similar motivation, our approach and problem formulation differ from theirs.
@@ 3.1 Problem Setting
Decision-making environment.
We define the decision-making environment as a Markov Decision Process (MDP), represented by $W=\langle S,A,P,R\rangle$. Here, $S$ denotes the set of states, $A$ represents the set of agent actions, $P(s^{\prime}|s,a)$ is the transition probability from state $s$ to state $s^{\prime}$ upon taking action $a\in A$, and $R(s,a,s^{\prime})$ represents the bounded reward received after taking action $a$ in state $s$ and reaching state $s^{\prime}$. To emphasize the goal recognition aspects of the problem, we introduce a set of goal states $G\subseteq S$. These goal states are terminal; that is, $P(g|g,a)=1$ $\forall g\in G,a\in A$.
Models of behavioral agents.
We represent the agent’s decision-making policy in a general form $\Pi:S\times T\rightarrow A$. Specifically, for an agent with a decision-making policy $\pi\in\Pi$, the agent will execute the action $\pi(s,t)$ in state $s$ at time $t$. The agent is conceptualized as a planner $H:W\rightarrow\Pi$, where the input is an environment $w\in W$, and the output is a policy $\pi\in\Pi$. To illustrate our formulation, consider a agent parameterized by a time-variant discounting function $d(t)$. The standard optimal agent model corresponds to a fixed discounting factor $\gamma\in(0,1]$ with $d(t)=\gamma^{t}$. We can also represent an agent with present bias (O’Donoghue and Rabin, 1999) by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$. We would like to highlight that our approach not only accounts for standard analytical closed-form expression of agent behavior. It can also accounts for scenarios where an agent policy $\pi$ is a machine learning model, i.e., a neural network trained on human behavioral data.
Worst-case distinctiveness (wcd).
To evaluate the difficulty of goal recognition, we adhere to standard literature and focus on the measure of worst-case distinctiveness (wcd(Keren et al., 2014), defined as the maximum number of steps an agent can take before revealing its goal. To compute the wcd for a given suboptimal agent $h\in H$ in an environment $w\in W$, we evaluate the path for the agent to each goal and compute the number of actions from the initial state that are identical for every goal.
@@ 3.3 Our Proposed Method
Predictive model for wcd.
To build the predictive model for wcd, we first curate a training dataset through simulations. For a given environment $w$ and agent behavioral model $h$, we can obtain $\textit{wcd}(w,h)$ by solving for the agent’s actions towards different goals. After collecting a training dataset, we train the predictive model using a convolutional neural network. Implementation details are included in the appendix.
Optimization procedure.
After obtaining the predictive model, we develop a gradient-based optimization framework. Note that this optimization framework strictly generalizes the existing literature in goal recognition design where the space of modifications are limited (e.g., usually limits to blocking an action in MDP) (Keren et al., 2021).
We then perform gradient descent on the relaxed Lagrangian. As environment modifications are often discrete (e.g., whether to block a cell in the grid world), we apply a discrete gradient descent procedure. Specifically, at each step of gradient descent, we obtain a gradient, which is a vector indicating the suggested change magnitude for each element (such as a cell in the grid world). We then select the element with the highest gradient value and make the corresponding change. Note that some suggested modifications may not be valid; for instance, we cannot block a cell that is already blocked in the grid world. In such cases, we proceed to the element with the next highest gradient value, continuing this process until a valid modification is made. This modification procedure is repeated until the gradient descent converges.
Note that with this Lagrangian relaxation, we cannot directly set the budget, however, based on duality, a larger Lagrangian multiplier $\lambda$ corresponds to a smaller budget $B$, in the original constrained optimization formulation. In our experiments, instead of selecting different budgets directly, we choose varying Lagrangian multipliers and record the realized costs of the modifications.
4 Experiment Setup
4.1 Benchmark Domains
We utilize two benchmark domains. The first is the standard basic grid world environment, commonly used in goal recognition design literature. The second is the Overcooked-AI environment (Carroll et al., 2019), an complex environment with a richer set of environment modification. This environment is particularly relevant to the downstream implications of our work, namely in supporting human-AI collaboration.
@@ 4.1.1 Benchmark Domain: Grid world
Design space of environment modifications.
In the context of goal recognition design in grid world, the space environment modification is often limited to adding blocks to spaces (Keren et al., 2014), also called action removal in the literature. In our work, we broaden the design space for modifications to also consider the unblocking of existing blocked spaces as potential modifications.
@@ 4.1.2 Benchmark Domain: Overcooked-AI
We also conduct our evaluations on a more complex domain: Overcooked-AI [^1] https://github.com/HumanCompatibleAI/overcooked_ai. This environment is based on the popular multiplayer game Overcooked. In each game, players (agents) collaborate to prepare and deliver specific recipes. Since the goal of our approach is to enable efficient inference of agent goals, we focus on the special case with a single agent. The environment is represented as a grid (see Figure 2(b)) with each cell specifying the object that is placed at the cell. The objects may be a counter, a tomato, an onion, a pot, a dish, a serving point, or an empty space. The agent can only occupy empty spaces in the environment and cannot step on any other objects in the environment. The agent can carry the movable objects, i.e. onion, tomato, and dish, and drop them on any of the other non-space objects (counter, serving point, or pot). To navigate in the environment, the agent can move left, right, north, or south and maintain an orientation that is consistent with the last movement direction.
Goal recognition in Overcooked-AI.
The goal of the agent in Overcooked-AI is defined by the recipe it needs to complete. For instance, to prepare a soup with one onion and two tomatoes, the agent must pick up the ingredients, place them in the pot, and cook them. In the context of goal recognition, the objective is to deduce which recipe the agent is preparing. For the sake of simplicity in our experiments, we focus on scenarios with only two possible goals or recipes.
Design space of environment modifications.
A primary challenge in the Overcooked-AI environment is its considerably larger design space for environment modifications. The design space includes the addition, removal, or changing position of any object within the environment, subject to the modification being valid. This means an object cannot be added to a cell already occupied by another object, and an object that does not exist cannot be removed. The objective of conducting experiments in this domain is to assess whether our approach can be used to address more complex domains.

### v2

@@ Abstract
Goal recognition design aims to make limited modifications to decision-making environments with the goal of making it easier to infer the goals of agents acting within those environments. Although various research efforts have been made in goal recognition design, existing approaches are computationally demanding and often assume that agents are (near-)optimal in their decision-making. To address these limitations, we introduce a data-driven approach to goal recognition design that can account for agents with general behavioral models. Following existing literature, we use worst-case distinctiveness (wcd) as a measure of the difficulty in inferring the goal of an agent in a decision-making environment. Our approach begins by training a machine learning model to predict the wcd for a given environment and the agent behavior model. We then propose a gradient-based optimization framework that accommodates various constraints to optimize decision-making environments for enhanced goal recognition. Through extensive simulations, we demonstrate that our approach outperforms existing methods in reducing wcd and enhancing runtime efficiency in conventional setup. Morevoer, our approach also adapts to settings in which existing approaches do not apply, such as those involving flexible budget constraints, more complex environments, and suboptimal agent behavior. Finally, we have conducted human-subject experiments which confirm that our method can create environments that facilitate efficient goal recognition from real-world human decision-makers.
@@ 1 Introduction
With the rapid advancement of artificial intelligence (AI), there has been a surge in interest in human-AI collaboration, aiming to synergize human and AI capabilities across various domains such as gaming, e-commerce, healthcare, and workflow productivity. Designing AI agents to work alongside humans requires these agents to understand and infer human goals and intentions. While there has been abundant research in goal recognition [23, 24] that aims to infer human goals by observing their actions, this work focuses on the goal recognition design problem [12], where one needs to identify how to modify decision-making environments to enable better goal recognition.
@@
Our work is built on the goal recognition design problem formulated by Keren et al. [12]. They proposed the worst-case distinctiveness (wcd) metric, defined as the maximum number of decisions an agent can make without revealing its goal, to measure the difficulty of inferring the agent’s goal. They then aimed to modify the decision-making environment, through removing allowable actions from states, to optimize this measure. Since the introduction of this work, there have been several follow-up works that extend the formulation to deal with different settings, such as stochastic settings [28, 29] and partial observability [14]. More discussion can be found in the survey by Keren et al. [15].
@@
While there has been significant progress in goal recognition design, there are two main limitations in the literature. First, current approaches require evaluating the difficulty of goal recognition, i.e., worst-case distinctiveness (wcd), for a large number of potential modifications to the decision-making environment. Since each evaluation of wcd requires solving the optimal policy multiple times to each of the goals, this process is computationally demanding and limits the scalability of these approaches, especially in virtual domains where the number of possible environment configurations is large. Second, decision-making agents in goal recognition design are often assumed to be optimal. While there are efforts addressing suboptimal agents [13, 27], these primarily focus on settings where there is a limited number of deviations from optimal decision-making. This assumption of optimality is limiting in scenarios involving human agents, who are known to often systematically deviate from optimal decision-making due to cognitive and informational constraints [10, 19].
@@
To address these limitations, we propose a framework for data-driven goal recognition design with general agent behavior. To relax the optimality assumption, we explicitly incorporate models of agent behavior into the optimization framework to better represent behavioral agents. To tackle the computational challenges, our approach leverages data-driven methods for goal recognition design. The core idea involves building a machine learning oracle that predicts the difficulty of goal recognition (e.g., wcd) given a decision-making environment and an agent’s behavioral model. This oracle is trained on datasets generated from simulations, allowing for the evaluation of wcd in any given environment and agent model. Such an approach significantly accelerates the evaluation of wcd during run-time optimization. Once the machine learning oracle is established, we apply the general gradient-based optimization approach to minimize wcd, employing Lagrangian relaxation to manage constraints on environment modifications. This allows us to address more general forms of objectives and constraints that existing approaches in the literature cannot address.
@@
To evaluate our framework, we conducted a series of simulations and human-subject experiments. We start with simulations in the standard setup, with the grid world environment and optimal agent assumption. We show that our approach outperforms existing baselines in reducing worst-case distinctiveness (wcd) and demonstrates considerably better run-time efficiency. We then conducted additional simulations to showcase that our approach can generalize to settings beyond the capabilities of existing approaches in the literature. These include scenarios involving flexible budget constraints, more complex environments, and suboptimal agent behavior. Lastly, we have conducted human-subject studies demonstrating that our method can be leveraged to design environments that facilitate efficient goal recognition from real-world human decision-makers. The results highlight the potential of our approach to enable more efficient human-AI collaboration.
Contributions. The main contributions of this work can be summarized as follows.
@@
    We propose a data-driven optimization framework for goal recognition design. This framework comprises a predictive model that estimates the wcd for a given environment and a model of agent behavior. We then utilize a gradient-based optimization method for goal recognition design. The framework is flexible to accommodate various environments, design spaces, and agent behavior.
@@
    Through extensive simulations, we show that our framework not only outperforms existing approaches in goal recognition design within standard settings but also adapts to scenarios that existing approaches cannot handle, including general optimization criteria, complex environments, and suboptimal agent behavior.
@@
    Through human-subject experiments, we demonstrate that our approach can adapt to agent models trained on human behavior. Furthermore, our framework can create environments that enable more effective goal recognition with real-world humans. To the best of our knowledge, our work is the first to utilize human-subject experiments to evaluate the environment design for goal recognition.
@@ 2 Related Work
Our work contributes to the expanding field of human-AI collaboration. Recent research has indicated that optimizing AI alone is insufficient for maximizing the performance of human-AI teams [1, 2]. To develop truly collaborative AI agents, it is crucial to equip them with the ability to comprehend and predict the intentions and goals of their human counterparts. This challenge lies at the heart of goal recognition research [11, 23, 9, 24, 20, 17]. Our work specifically focuses on goal recognition design, an extension of goal recognition that includes modifying environments to better facilitate the process of recognizing goals.
Goal recognition design was formulated by Keren et al. [12]. Since this seminal work, numerous research efforts have expanded the concept to accommodate stochastic environments [29, 27], different levels of observability [14, 30], and a variety of design spaces [18]. The studies most closely aligned with our approach focus on suboptimal agents [13, 27]. However, these studies characterize suboptimality by limiting deviations from the optimal policy, a method that may not adequately represent behavior of human agents, who frequently deviate systematically from optimal decision-making. Additionally, most existing work requires evaluating the difficulty of goal recognition in run time for numerous environmental modifications, which constrains the scalability of these methods. Our work sets itself apart by broadening the goal recognition design question to incorporate general agent behavioral models and by implementing a data-driven optimization approach.
From a technical standpoint, we adopt a data-driven optimization approach for goal recognition design. The use of data-driven tools in optimization has gained increasing prominence in the field of mechanism design, as evidenced by a range of recent studies [5, 7, 8, 4, 16, 22, 21, 3, 32]. Moreover, our work closely aligns with the recent studies that explicitly encoded human behavior and responses in the design of computational and learning systems [25, 26, 31, 32, 6]. While our work shares a similar motivation, our approach and problem formulation differ from theirs.
@@ 3.1 Problem Setting
Decision-making environment. We define the decision-making environment as a Markov decision process (MDP), represented by $W=\langle S,A,P,R\rangle$. Here, $S$ denotes the set of states, $A$ represents the set of agent actions, $P(s^{\prime}|s,a)$ is the transition probability from state $s$ to state $s^{\prime}$ upon taking action $a\in A$, and $R(s,a,s^{\prime})$ represents the bounded reward received after taking action $a$ in state $s$ and reaching state $s^{\prime}$. To emphasize the goal recognition aspects of the problem, we introduce a set of goal states $G\subseteq S$. These goal states are terminal; that is, $P(g|g,a)=1$ $\forall g\in G,a\in A$.
Models of behavioral agents. We represent the agent’s decision-making policy in a general form $\Pi:S\times T\rightarrow A$. Specifically, for an agent with a decision-making policy $\pi\in\Pi$, the agent will execute the action $\pi(s,t)$ in state $s$ at time $t$. The agent is conceptualized as a planner $H:W\rightarrow\Pi$, where the input is an environment $w\in W$, and the output is a policy $\pi\in\Pi$. To illustrate our formulation, consider a agent parameterized by a time-variant discounting function $d(t)$. The standard optimal agent model corresponds to a fixed discounting factor $\gamma\in(0,1]$ with $d(t)=\gamma^{t}$. We can also represent an agent with present bias [19] by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$. We would like to highlight that our approach not only accounts for standard analytical closed-form expression of agent behavior. It can also accounts for scenarios where an agent policy $\pi$ is a machine learning model, i.e., a neural network trained on human behavioral data.
Worst-case distinctiveness (wcd). To evaluate the difficulty of goal recognition, we adhere to standard literature and focus on the measure of worst-case distinctiveness (wcd[12], defined as the maximum number of steps an agent can take before revealing its goal. To compute the wcd for a given suboptimal agent $h\in H$ in an environment $w\in W$, we evaluate the path for the agent to each goal and compute the number of actions from the initial state that are identical for every goal.
@@ 3.3 Our Proposed Method
The main benefits of our approach compared with existing methods include: 1) the ability to incorporate different agent models $h$, 2) flexibility for various forms of optimization objectives and constraints, and 3) run-time efficiency.
@@
Predictive model for wcd. To build the predictive model for wcd, we curate a training dataset through simulations. For a given environment $w$ and agent behavioral model $h$, we can obtain $\textit{wcd}(w,h)$ by solving for the agent’s actions towards different goals. After collecting a training dataset, we train the predictive model using a convolutional neural network. Implementation details are included in the appendix.
Optimization procedure. After obtaining the predictive model, we develop a gradient-based optimization framework. This framework strictly generalizes the existing literature in goal recognition design where the space of modifications are limited (e.g., usually limits to blocking an action in MDP) [15].
We then perform gradient descent on the relaxed Lagrangian. As environment modifications are often discrete (e.g., whether to block a cell in the grid world), we apply a discrete gradient descent procedure. Specifically, at each step of gradient descent, we obtain a gradient, which is a vector indicating the suggested change magnitude for each element (such as a cell in the grid world). We then select the element with the highest gradient value and make the corresponding change. Note that some suggested modifications may not be valid; for instance, we cannot block a cell that is already blocked in the grid world. In such cases, we proceed to the element with the next highest gradient value, continuing this process until a valid modification is made. This modification procedure is repeated until the gradient descent converges. Note that with this Lagrangian relaxation, we cannot directly set the budget, however, based on duality, a larger Lagrangian multiplier $\lambda$ corresponds to a smaller budget $B$, in the original constrained optimization formulation. In our experiments, instead of selecting different budgets directly, we choose varying Lagrangian multipliers and record the realized costs of the modifications.
4 Experiment Setup
4.1 Benchmark Domains
We utilize two benchmark domains. The first is the standard basic grid world environment, commonly used in goal recognition design literature. The second is the Overcooked-AI environment [2], an complex environment with a richer set of environment modification. This environment is particularly relevant to the downstream implications of our work, namely in supporting human-AI collaboration.
@@ 4.1.1 Benchmark Domain: Grid world
Design space of environment modifications. In the context of goal recognition design in grid world, the space environment modification is often limited to adding blocks to spaces [12], also called action removal in the literature. In our work, we broaden the design space for modifications to also consider the unblocking of existing blocked spaces as potential modifications.
@@ 4.1.2 Benchmark Domain: Overcooked-AI
We also conduct our evaluations on a more complex domain: Overcooked-AI [^1] https://github.com/HumanCompatibleAI/overcooked_ai. This environment is based on the popular game Overcooked. In each game, players (agents) collaborate to prepare and deliver specific recipes. Since the goal of our approach is to enable efficient inference of agent goals, we focus on the special case with a single agent. The environment is represented as a grid (see Figure 2(b)) with each cell specifying the object that is placed at the cell. The objects may be a counter, a tomato, an onion, a pot, a dish, a serving point, or an empty space. The agent can only occupy empty spaces in the environment and cannot step on any other objects in the environment. The agent can carry the movable objects, i.e. onion, tomato, and dish, and drop them on any of the other non-space objects (counter, serving point, or pot). To navigate in the environment, the agent can move left, right, north, or south and maintain an orientation that is consistent with the last movement direction.
Goal recognition in Overcooked-AI. The goal of the agent in Overcooked-AI is defined by the recipe it needs to complete. For instance, to prepare a soup with one onion and two tomatoes, the agent must pick up the ingredients, place them in the pot, and cook them. In the context of goal recognition, the objective is to deduce which recipe the agent is preparing. For the sake of simplicity in our experiments, we focus on scenarios with only two possible goals or recipes.
Design space of environment modifications. A primary challenge in the Overcooked-AI environment is its considerably larger design space for environment modifications. The design space includes the changing position of any object within the environment, subject to the modification being valid. This means that the change cannot result in a new design where any of the objects is unreachable by the agent. The objective of conducting experiments in this domain is to assess whether our approach can be used to address more complex domains.